OpenStudy (angelwings996):

Simplify the difference 3y+1/4y+4 - 2y+7/2y+2

OpenStudy (angelwings996):

\[\frac{ 3y+1 }{ 4y+4 } - \frac{ 2y+7 }{ 2y+2 }\]

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ 3y+1 }{ 4y+4 } - \frac{ 2y+7 }{ 2y+2 }\] \[\Large \frac{ 3y+1 }{ 4(y+1) } - \frac{ 2y+7 }{ 2(y+1) }\] \[\Large \frac{ 3y+1 }{ 4(y+1) } - \frac{ 2(2y+7) }{ 2*2(y+1) }\] \[\Large \frac{ 3y+1 }{ 4(y+1) } - \frac{ 4y+14 }{ 4(y+1) }\] Do you see what to do from here?

jimthompson5910 (jim_thompson5910):

The first goal is to get every denominator to be the same number or expression (equal to the LCD) Once the denominators are equal, you can add or subtract the fractions

OpenStudy (mathstudent55):